The Response of Escherichia coli Growth Rate to Osmotic Shock

Enrique R. Rojas^{1,2}, Julie Theriot^{2,3,*,a}, and Kerwyn C. Huang^{1,3,*,a}

1Department of Bioengineering, Stanford University, Stanford, California,

USA 94305

2Department of Biochemistry and Howard Hughes Medical Institute, Stanford

University School of Medicine, Stanford, California 94305, US 3Department of Microbiology and Immunology, Stanford University School of

Medicine, Stanford, California 94305, USA

*To whom correspondence should be addressed E-mail: theriot@stanford.edu, kchuang@stanford.edu

aThese authors contributed equally to this work.

Abstract

Several recent theoretical studies have proposed that osmotic pressure plays an essential role during bacterial growth, positing that it is required to stretch the cell wall in order for the cell to elongate. To test this theory we monitored single-cell elongation rate of Escherichia coli while changing the osmolarity of its growth medium. We find several main results. First, although the steady-state elongation rate of E. coli is negatively correlated with osmolarity of the medium, after hyperosmotic shock the bacterium is able to maintain its original elongation rate for a period of tens of minutes. Second, oscillatory hyperosmotic shock reveals that temporary plasmolysis can slow cell elongation, but that the bacterium can “store” growth, such that when osmotic pressure is re-established the cells elongate to the length that they would have attained had they never experienced the shock. Third, MreB motion is unaffected by hyperosmotic shock. Finally, in response to hypoosmotic shock E. coli attains its steady-state elongation rate over the course of ≈5 min, and mutant strains lacking stretch-activated ion channels actually elongate slower after hyperosmotic shock than do wild-type E. coli. Furthermore, none of these bacteria undergo signficant osmotic swelling during hypoosmotic shock. These results lead us to present a model of cell wall expansion in which osmotic pressure plays only a minor role and that peptidoglycan synthesis is the sole downstream rate-limiting step in bacterial growth.

Cell growth is a complex process that integrates biochemical processes with cell-scale mechanical forces. For plant, fungal, and bacterial cells, growth requires both the accumulation of cytoplasm and the expansion of the cell wall, the extracellular polymeric network that defines the shape of these cells. In some cases, the biomechanical factors that mediate this expansion are known. For example, osmotic pressure drives gradual deformation of the cell wall during plant cell growth [1, 2]. However, the bacterial cell wall is distinct from its eukaryotic counterparts, in both ultrastructure and chemical composition. In particular, the cell wall of Gram-negative bacteria, which is composed entirely of peptidoglycan, is perhaps as thin as a molecular monolayer [3]. Therefore, while several theoretical studies have proposed that bacterial cell wall expansion is also pressure-mediated [4, 5, 6, 7], this theory needs to be tested experimentally.

Osmotic pressure is the pressure that forces water into or out of a cell due to a differential in osmolarity across the plasma membrane, and is given by the Morse Equation, P = RT (C_{in} − C_{out}), where C_{in} is the osmolarity of cytoplasm, C_{out} is the osmolarity of the extracellular medium, R is the gas constant, and T is the temperature. Two experimental methods have been used to measure P in the Gram-negative bacterium Escherichia coli and yielded values of 1-3 atm [8, 9]. A primary role of the cell wall is to prevent lysis by balancing this load with mechanical stress. As early as 1924, Walter suggested that this stress, in turn, may be exploited by the bacteria. He set forth a theory of bacterial growth based on osmotic swelling of the cell and concomitant stress relaxation within the cell wall [10]. In support of this theory, he and others found that the growth rate of a number of bacterial species, including E. coli, is inversely correlated with the osmolarity of the growth medium ([11, 12, 13], Fig. 1A). They also showed that this effect does not depend on the chemical used to modulate the osmolarity, demonstrating that the cause of growth-rate decrease is not due to specific, toxic reactions, but is essentially osmotic. However, since bacteria possess a number of mechanisms to regulate their cytoplasmic osmolarity in response to changes in their osmotic environment [14], it is unclear whether raising medium osmolarity causes a decrease in osmotic pressure, as would be expected from the Morse Equation. This inverse correlation between C_{out} and P could be the result of another osmotic effect such as altered water potential within the cell (which could affect protein folding [15], signaling [16], or other colligative properties of the cytoplasm), or could result from a combination of such effects (Fig. 1B).

Notwithstanding, several theorists have built upon Walter’s seminal idea to account for information regarding the molecular composition of the bacterial cell wall. Koch et al. proposed a model in which cell wall expansion and commensurate cell growth occur when load-bearing cross-links within the wall are selectively hydrolyzed and the network stretches incrementally to balance the osmotic pressure [4]. Recently, Furchtgott et al. simulated expansion of the Gram-negative cell wall by treating it as a dynamic network of linear springs under tension [5]. Although pressure-driven wall expansion is essential to these models, neither was used to quantitatively predict the scaling between osmotic pressure and growth rate. However, Jiang and Sun derived a mean-field model of bacterial growth by balancing the mechanical energy of cell wall deformation with the chemical energy of wall synthesis. This model predicts a Figure 1: Escherichia coli maintains its elongation rate after hyperosmotic shock. (A) The population-averaged steady-state elongation rate of E. coli as a function of the concentration of sorbitol in the growth medium (◯). Also shown is the population-averaged elongation rate after recovery from 400mM hyperosmotic shock (i.e., during Phase III, △) and the populationaveraged elongation rate immediately after 400mM hypoosmotic shock (□). (Inset) Two possible pathways by which extracellular osmolarity could modulate elongation rate. (Outset) Flow chart describing the temporal response of elongation rate to osmotic shock. Error bars indicate ±1 s.e. (B) E. coli stained with WGA, imaged in both phase and epifluorescence. (C) Kymographs of a cell elongating during 400mM hyperosmotic shock. The black arrow indicates the formation of a septum during cell division. The white arrow indicates the period after the shock during which the cytoplasm is shorter than the cell wall. (D) The concentration of sorbitol in the growth medium versus time during a hyperosmotic shock. The osmolality of the LB growth medium alone is 260 mMol/Kg. (E) The length of the cell wall of representative cells during the same shock. (F) The population-averaged elongation rate versus time of both the cytoplasm and the cell wall during the shock showing the four phases of response. quadratic dependence between osmotic pressure and the elongation rate of rod-shaped bacteria [6]. A novel role for osmotic pressure was proposed by Amir and Nelson, who formulated a model of cell-wall expansion based on propagating dislocations in the peptidoglycan network [7]. According to this model, the rate of cell growth is proportional to the speed at which the protein complexes that synthesize peptidoglycan move within the cell, which is, in turn, proportional to the osmotic pressure.

Indeed, all of the above models implicitly rely on both synthesis of new peptidoglycan and turnover of old peptidoglycan to achieve cell wall expansion during growth. A popular opinion is that the biosynthetic and hydrolytic enzymes form a complex, thereby ensuring spatial and temporal coordination of the two processes, but this idea is purely speculative [17]. However, three D,D-endopeptidases, which hydrolyze peptide bonds within the peptidoglycan, were recently identified in E. coli. As a group, these enzymes were shown to be essential for incorporation of nascent peptidoglycan into the cell wall, and for cell growth [18], which suggests a degree of coupling between synthesis and hydrolysis of peptidoglycan.

The relationships between cell wall turnover, cell wall synthesis, the dynamics of the synthesis machinery, and mechanical stress likely constitute a system that determines the rate of cell wall expansion, as well as the response of bacterial growth rate to changes in osmotic pressure. Given the abundance of theory and speculation regarding these relationships, we altered the osmotic pressure of E. coli via osmotic shock and measured the effect that this had on the single-cell elongation growth rate, and on the dynamics of MreB, a protein whose motion is dependent on peptidoglycan synthesis [19, 20, 21]. By examining these responses across time scales that span three orders of magnitude, we were able distinguish the effect of osmotic pressure from that of medium osmolarity on cell growth. In contrast to the predictions of the models described above, we found that growth rate is insensitive to moderate changes in osmotic pressure. Thus, our results rule out a mechanically based model of cell wall expansion in E. coli, and indicate that the rate of cell wall synthesis is the sole, rate-limiting step in cell growth.

Results

E. coli maintains its elongation rate after hyperosmotic shock

Because E. coli cells elongate exponentially [22], their growth rate can be expressed as a relative rate of elongation, ė = \frac{1}{l} \frac{dl}{dt}, where l is the axial length of the cell. This empirical quantity can account for both reversible elongation of the cell, in which the cytoplasm acquires water but where the cell wall stretches elastically, and irreversible elongation of the cell, in which uptake of water is accompanied by hydrolysis and/or synthesis of peptidoglycan. While the former type of elongation is likely to occur in response to osmotic shock, the latter type occurs during steady-state elongation and may also have a pressure dependence.

We measured the steady-state elongation rate of single E. coli cells as a function of medium osmolarity and found a negative correlation between these quantities (Fig. 1A), in agreement with the classic bulk measurements [11, 12, 13]. To distinguish between the effect of osmotic pressure versus that of osmolarity alone, we measured the time scale at which elongation rate relaxes to its steady-state value in response to changes in medium osmolarity. We reasoned that if osmotic pressure drives elongation, then upon hyperosmotic shock, which dehydrates the cell within seconds (Fig. S1, [23]), the elongation rate would quickly attain its steady-state value at the new osmolarity, or would fall below the steady-state value until solute transporters upregulate osmotic pressure to its steady-state value.

To examine this idea, we fluorescently labeled the cell wall and took time-lapse images of E. coli with phase and epifluorescence microscopy (Fig. 1C,D) during 400mM hyperosmotic shocks (Fig. 1 D,E,F). There are four stages of response to the hyperosmotic shock (Fig. 1F,G). I) First, during the shock, both the cytoplasm and the cell wall predictably shrink due to dehydration of the cell and consequent elastic contraction of the cell wall (Fig. 1D,E). The cytoplasm shrinks significantly more than the cell wall, i.e., the cell plasmolyzes (Fig. 1D). Separation of the cell wall and the cytoplasm is most commonly seen at the pole(s) of the cell. II) During the following several minutes, the cytoplasm rapidly elongates while the cell wall elongates at a slower rate (Fig. 1G). The time scale of recovery from plasmolysis is in agreement with a recent study that examined the effect of larger hyperosmotic shocks on E.

Figure 2: Hyperosmotic shock causes stored growth. (A) The concentration of sorbitol in the growth medium during 100 mM oscillatory hyperosmotic shock with 90 s period. (B) The population-averaged width of the cell wall during the shock. The confidence intervals indicate ±1 s.e. (C) The length of the cell walls of a number of representative cells during the shock. (D) The population-averaged elongation rate of the cell wall during the shock. (E) The populationaveraged elongation rate of both the cytoplasm and the cell wall versus phase of the oscillatory cycle. Error bars indicate ±1 s.e. (F) “Population-averaged” length of the cytoplasm and cell wall obtained by integrating their average elongation rate. (Inset) Zoom-in of two cycles.

coli cell volume [23]. III) During the 10-40 minutes after recovery from plasmolysis the cells elongate at a rate approximately equal to the pre-shock elongation rate, and therefore faster than their steady-state value at the new osmolarity (Fig 1A,G). IV) Finally, the bacteria decelerate to their steady-state value. This general behavior is independent of the pre-shock osmolarity (Fig. 1A,S2).

The fact that cells resume their pre-shock elongation rate after recovering from plasmolysis, during which the turgor pressure (the normal force per unit area exerted by the cytoplasm on the cell wall) is necessarily zero, suggests that elongation rate is not regulated by mechanical stress within the cell wall. Whereas osmotic pressure changes immediately upon osmotic shock, elongation rate appears to be a function of variables that vary slowly, on the order of tens of minutes after the shock. Strikingly, we observe a significant rate of cell wall elongation even during plasmolysis (Fig. 1G, Phase II), which demonstrates that mechanical stress is not even required for cell wall elongation. This basal elongation may be due to cell wall synthesis in the absence of turgor pressure.

E. coli stores growth during oscillatory hyperosmotic shock

From these single-shock experiments alone we could not rule out the possibility that osmotic pressure re-equilibrates or even overshoots its initial value within minutes after hyperosmotic shock, and that this accounts for the resumption of the initial elongation rate. E. coli responds to hyperosmotic shock by importing potassium, as well as compatible solutes such as proline and glycine betaine, if they are present [24]. Potassium takes several minutes to accumulate in the cytoplasm [25], while the time for activation of the broad-specificity compatible solute transporter ProP is 1-1.7 min [26]. Also, in response to large hyperosmotic shock, E. coli recovers from plasmolysis within several minutes (Fig. 1G, [23]), indicating this may be the time scale of osmoregulation. Therefore, we wondered if pressure-dependent effects on elongation rate would be evident at shorter time scales, < 1 min, after hyperosmotic shock, before osmoregulatory mechanisms are likely to be able to achieve osmotic homeostasis.

It is difficult to measure elongation rates at sub-minute time scales due to the size of bacteria: in Luria broth (LB), E. coli cells elongate at ≈ 2 hr^{-1} (Fig. 1A), which equates to ≈2 nm/s for a typical cell. In order to precisely measure elongation rate on short time scales, in spite of the optical diffraction limit, we combined three strategies: 1) we determined the cell outlines to sub-pixel resolution and then were able to discern changes in this location below the diffraction limit; 2) we averaged the elongation rate over a population of cells; and 3) we averaged the elongation-rate response of these populations over many successive identical osmotic shocks. To achieve the latter, we subjected the bacteria to oscillatory hyperosmotic shock, during which, after an initial period in LB, the medium was repeatedly exchanged for LB concentrated with sorbitol (Fig. 2A, Movie S1). Finally, to avoid significant plasmolysis, we decided to examine the effects of relatively small-amplitude osmotic shocks, <300 mM. Across this range, the variation in steady-state elongation rate is small or negligible (Fig. 1A), which allows us to isolate the effect of the change in osmotic pressure versus that of the osmolarity.

The elongation rate of E. coli during oscillatory hyperosmotic shock depends on the phase of the oscillatory cycle. First, as the cells elongate, they acutely shrink and swell, due to osmosis, during the phases of the cycle when the medium osmolarity is being raised or lowered (Fig. 2B,C). The cell-wall width decreases upon each hyperosmotic shock, and this decrease is Figure 3: Hyperosmotic shock does not affect MreB motion. (A) 100 second-long average intensity projections of E. coli expressing a MreB-GFP fusion protein. (B) The concentration of sorbitol in the growth medium and the average velocity of MreB puncta during oscillatory hyperosmotic shock. (C) The average velocity of MreB puncta versus phase of the oscillatory cycle. Error bars indicate ±1 s.e.

reversed only when the bacteria are re-immersed in LB, suggesting that we are indeed reducing osmotic pressure for the duration of each shock (Fig. 2B). The cell-wall length also undergoes acute deformation during the phases of media exchange, which leads to sharp peaks and valleys in the the time-dependent cell-wall elongation rate (Fig 2D).

During the phases of the cycle when the cells are in LB, the their cell walls elongate precisely as fast as they did before the oscillatory shock, while in the same medium (Fig. 2E). Conversely, during the phases of concentrated media, the cell wall elongates at a significantly slower rate. Even though the bacteria spend equal time in each media, the cell-wall elongation rate averaged over the course of the entire oscillatory shock is not equal to the average of the elongation rates in the two media, but is equal to the steady-state elongation rate in LB alone. This paradoxical result is resolved by the fact that the cell-wall elongation rate, averaged over the phases of media exchange, is higher than the steady-state elongation rate in LB.

In contrast to that of the cell wall, the elongation rate of the cytoplasm (Fig. S3) does not show a phase dependence and is, in fact, equal to the steady-state elongation rate in LB throughout the oscillatory hyperosmotic shock (Fig. 2E). This fact suggests that the cell wall and the cytoplasm can move independently during these experiments. To examine this, we obtained “population-averaged” time-dependent lengths of the cell wall and the cytoplasm by integrating their average elongation-rate traces, assuming a typical initial length of 3 μm (Fig.

2F). These traces reveal that the cytoplasm shrinks slightly more than the cell wall upon each hyperosmotic shock, and proceeds to elongate faster than it until the length of the two are equal. Thus, as in the single-shock experiments, the elongation of the cell wall is only retarded when the cell is plasmolyzed. Since the separation of the cell wall from the cytoplasm is 10-20 nm we did not resolve this slight plasmolysis in single cells, which demonstrates the power of averaging the elongation rate over a population of cells.

The above behavior was consistent over the entire range of oscillation periods that we were able to apply (>60 s, Fig. S4) and across a range of shock magnitudes (0-100mM). As a whole, these data suggest that the cell-wall elongation rate is pressure-independent, unless the cell is plasmolyzed, i.e. P ≤ 0. However, we propose that even during slight plasmolysis, in which the inner membrane recedes from the cell wall at the pole(s) of the cell, the cells continue to incorporate peptidoglycan into the lateral cell wall, albeit in a unextended state. This explains why there is a non-zero cell-wall elongation rate during plasmolysis, and also why when pressure is re-established the cell wall extends to the length that it would have attained had pressure never been depleted in the first place. This behavior is reminiscent of so-called “stored growth,” long observed in certain plant tissues [27] and more recently in unicellular algae [28].

For shock magnitudes between 100-300mM, cells undergo more severe, visible plasmolysis during each cycle. Although they are still able to store growth, their average elongation rate during the oscillatory shock is lower than the steady-state rate in LB (Fig. S4). To explain this, we hypothesize that during severe plasmolysis, the the inner membrane recedes from the lateral cell wall as well as the poles and the cell wall is no longer able to incorporate nascent peptidoglycan during the entire cycle.

Hyperosmotic shock does not affect MreB motion

The previous interpretation rests on the assumption that the rate of addition of cell-wall material is unaffected by hyperosmotic shock, despite the decrease in elongation rate. To infer the rate of cell wall addition, we examined the dynamics of MreB (Fig. 3A), a protein required for rod morphology and whose circumferential motion is dependent on cell-wall synthesis in E.

Figure 4: Elongation rate equilibrates slowly after hyposmotic shock. (A) The populationaveraged elongation rate of three strains of E. coli during hypoosmotic shock: 1) MG1655 (wild type), 2) MJF465 (lacking mscS and mscL), and 3) MJF612 (lacking mscS, mscL,mscM, and mscK). (B) Comparison of length of representative cells during 400 mM hypoosmotic shock and 400mM hyperosmotic shock.

coli [19]. When cells expressing an MreB-GFP fusion protein were subjected to oscillatory hyperosmotic shock, the average speed of MreB puncta remained constant, even during phases when the elongation rate of the cell wall was decreased (Fig 3B,C). These data provide further evidence that, although it is possible to transiently slow E. coli elongation via plasmolysis, cell wall synthesis ultimately determines cell length.

Hypoosmotic shock does not cause swelling of E. coli

We next inquired whether cell wall synthesis is also rate-limiting under hypoosmotic shock (Fig. 4A). As with hyperosmotic shock, E. coli does not immediately obtain its steady-state elongation rate after hypoosmotic shock, but gradually accelerates to it over the course of ≈ 5 min (Fig. 4A,1A), indicating that this acceleration is not pressure-mediated.

Interestingly, we observe virtually no swelling of the cell during hypoosmotic shock (Fig. 4B). The lone evidence of osmotic swelling is a small peak in elongation rate during the shock (Fig. 4A). This is in sharp contrast to the drastic shrinking that we see upon hyperosmotic shocks of the same magnitude, and to the swelling that we see during the periodic downshocks that occur as part of the oscillatory hyperosmotic shock (Fig. 2C).

One possible explanation for these results is that down-regulation of osmotic pressure is faster than the time of media exchange, and therefore it is essentially impossible to increase osmotic pressure beyond its steady-state value. Indeed, stretch-activated ion channels, which release cytoplasmic solutes passively, are observed, in vitro, to open within fractions of a second in response to mechanical stress in the plasma membrane [29], and their opening probability is Figure 5: Conceptual model of stored growth in E. coli. Green bars represent glycan polymers while red bars represent polypeptides and the blue sphere represents peptidoglycan biosynthetic apparatus, which includes MreB. New glycan polymers are inserted such that cross-linking polypeptides are looped out of the plane, and are subsequently extended if osmotic pressure is positive. During periods of plasmolysis, elongation is “stored” as polypeptide loops, that are fully extended when pressure is restored. found to be switch-like around a threshold value [30]. To test the contribution of these channels to the growth-rate response of E. coli to hypoosmotic shock, we used strains of E. coli that lacked two or more of the genes encoding the seven known channels. Not even bacteria that lacked the channels MscS and MscL showed measurable swelling during hypoosmotic shock (Fig. 4B), and the lack of these genes actually reduced the elongation rate after the shock as compared with the wild type bacteria (Fig. 4A). 99% of cells possessing the MscS-MscL double knockout continued to elongate after 400mM hypoosmotic shock. On the other hand, 0% of E. coli that lack four of the ion channels (MscS, MscL, MscM, MscK) continued to elongate after such a shock (Fig. 4C,D), and many of these cells underwent visible cytolysis. This survival rate is consistent with a recent study concerning the contributions of the various ion channels to the ability of E. coli to cope with hypoosmotic shock [31]. Evidently, it is impossible to increase the length of E. coli cells beyond the length prescribed by cell wall synthesis without causing lysis. This result further supports the idea that synthesis is the ultimate rate-limiting step during cell elongation.

Discussion and Conclusions

In order to measure the fundamental dependence of E. coli growth on osmotic pressure we monitored single-cell elongation rate and MreB dynamics during osmotic shock. This is the first such study in bacteria. We conclude that osmotic pressure is not an essential driving force of cell wall expansion in this organism.

First, the elongation rate of E. coli responds on slow time scales (>5 min) to both hyperand hypo-osmotic shock. Second, osmotic shock has no effect on MreB motion. Third, although is it possible to slow cell wall expansion by plasmolyzing the cell (i.e., reducing the osmotic pressure below zero), the bacteria can “store” growth such that when osmotic pressure is re-established the cell elongates to the length that it would have attained had it never been plasmolyzed. Finally, hypoosmotic shock cannot extend the cell wall longer than the length determined by steady-state elongation.

As a whole, these data are consistent with a model in which the rate of cell wall synthesis, which is independent of osmotic pressure, is the primary determinant of cell elongation rate. To account for stored growth, we propose that nascent peptidoglycan is inserted into the cell wall in an “unextended,” state, and that positive osmotic pressure is required to extend it (Fig. 5). Thus, during plasmolysis, although the rate of insertion is unaffected, this material makes a smaller contribution to elongation than when the cell is turgid. However, if pressure is reestablished, this material then contributes its full potential to cell elongation. That is, according to this model, osmotic pressure does play a trivial role whereby a non-zero value is required to stretch out newly synthesized, un-extended peptidoglycan. This model is a simple elaboration of the long-standing “make-before-break” hypothesis, in which elongation is exactly equal to insertion of peptidoglycan subunits [32], and our data constitute the first evidence in support of this model.

Our data contradict several theoretical studies that predict that bacterial elongation rate and/or MreB velocity scales directly as osmotic pressure [6, 7], and provide a fundamental constraint with which to test other other mechanical models [4, 5], including future ones. Because elongation rate does not respond rapidly to changes in the osmolarity of the growth medium, it is unlikely that the dependence between cell elongation rate and medium osmolarity [10, 11, 12, 13] is mediated by osmotic pressure, although we can not rule out a scenario in which osmotic pressure is involved in a signaling cascade that modulates elongation rate.

The data from E. coli also contrast the wealth of evidence from plants that show that cell elongation in these organisms is directly tunable by osmotic pressure. On the other hand, E. coli share with plants the capability to store growth upon depletion of turgor pressure [27, 28]. In the case of E. coli, which has a very thin (≤3 nm, [3]) cell wall, it is easy to imagine a mechanism for stored growth that depends on “coiled insertion,” as described above (Fig. 5), whereas in plants, which have a much thicker (>100 nm, [33]) cell wall, the mechanisms of storing growth are more speculative.

A surprising finding of our study was that upon hypoosmotic shock, E. coli shows virtually no osmotic swelling or elastic expansion of the cell wall. In fact, not even bacteria that lyse in response to the shock show swelling during lysis. This suggests that the cell wall may be storing stress without stretching, and hints at a second possible reason why there is no elastic expansion upon hypoosmotic shock even though there is significant contraction upon hyperosmotic shock: that the cell wall has non-linear mechanical properties. Indeed, the E. coli cell wall has been shown to strain-stiffen in response to AFM indentation [34]. In this light, it is important to note that although the mechanical integrity of the E. coli cell is commonly attributed to the cell wall, it is also possible that the outer membrane, which is rich in cross-linked lipopolysaccarides, is also load-bearing, and the mechanical properties of the composite cell envelope requires a theory more sophisticated than linear elasticity.

E. coli growth is evidently robust to changes in osmotic pressure. This could be an important adaptation for an enteric bacterium, which may face drastically different osmotic environments during entry and exit from the host. It is necessary to perform similar studies on other species of bacteria to see if the behavior we observe for E. coli is universal. In particular, it will be interesting to examine Gram-positive bacteria, which have a thicker cell wall (≈40 nm, [35]), to see whether this feature is sufficient to endow them with a more plant-like elongation-rate response to changes in osmotic pressure.

Acknowledgments

Funding for this project was provided by an NIH New Innovators Award and by the Howard Hughes Medical Institute. E.R. was supported by a postdoctoral fellowship from Simbios, an NIH-funded center for computational biophysics at Stanford University.

References

[1] Green PB (1968) Growth physics in Nitella: a method for continuous in vivo analysis of extensibility based on a micro-manometer technique for turgor pressure. Plant Physiol. 43: 1169-1184.

[2] Proseus TE, Zhu G-L, Boyer JS (2000) Turgor, temperature and the growth of plant cells: using Chara corallina as a model system. J Expt Bot. 51: 1481-1494.

[3] Gan L, Chen S, Jensen, GJ (2008) Molecular organization of Gram-negative peptidoglycan. Proc Natl Acad Sci USA. 105:18953-18957.

[4] Koch AL, Higgins ML, Doyle RJ (1982) The role of surface stress in the morphology of microbes. J Gen Micro. 128:927-945.

[5] Furchtgott L, Wingreen NS, Huang KC (2011) Mechanisms for maintaining cell shape in rod-shaped Gram-negative bacteria. Mol Microbiol. 81:340-353

[6] Jiang H, Sun SX (2010) Morphology, growth, and size limit of bacterial cells. Phys Rev Lett. 105:028101.

[7] Amir A, Nelson DR (2012) Dislocation-mediated growth of bacterial cell walls. Proc Natl Acad Sci USA. 109:9833-9838.

[8] Cayley DS, Guttman HJ, Record TR (2000) Biophysical characterization of changes in amounts and activity of Escherichia coli cell and compartment water and turgor pressure in response to osmotic stress. Biophys J. 78:1748-1764.

[9] Deng Y, Sun M, Shaevitz JW (2011) Direct measurement of cell wall stress stiffening and turgor pressure. Phys Rev Lett. 107:158101.

[10] Walter H (1924) Plasmaquellung und wachstum. Z Bot. 16:353-417.

[11] Scott WJ (1953) Water relations of Staphylococcus aureus at 30^{∘}C. Aus J Biol Sci. 6:549-564

[12] Christian JHB, Scott WJ (1953) Water relations of Salmonellae at 30^{∘}C. Aus J Biol Sci. 6:565-573

[13] Christian JHB (1955) The influence of nutrition on the water relations of Salmonella oranienburg. Aus J Biol Sci. 8:75-82.

[14] Wood JM (2006) Osmosensing by bacteria. Sci STKE. 357:pe43.

[15] Levy Y, Onuchic JN (2006) Water mediation in protein folding and molecular recognition. Annu Rev Biophys Biomol Struct. 35:389-415.

[16] Miermont A, et al. (2013) Severe osmotic compression triggers a slowdown of intracellular signaling, which can be explained by molecular crowding. Proc Natl Acad Sci USA. 110:5725-5730.

[17] Höltje JV (1998) Growth of the stress-bearing and shape-maintaining murein sacculus of Escherichia coli. Microbiol Mol Biol Rev. 62:181203.

[18] Singh SK, SaiSree L, Amrutha RN, Reddy M (2012) Three redundant murein endopeptidases catalyse an essential cleavage step in peptidoglycan synthesis of Escherichia coli K12. Mol Microbiol. 86:1036-1051.

[19] van Teeffelen S, et al. (2011) The bacterial actin MreB rotates and rotation depends on cell-wall assembly. Proc Natl Acad Sci USA. 108:15822-15827.

[20] Garner EC, et al. (2011) Coupled, circumferential motion of the cell wall synthesis machinery and MreB filaments in B. subtilis. Science. 333:222-225.

[21] Domínguez-Escobar J, et al. (2011) Processive movement of MreB-associated cell wall biosynthetic complexes in bacteria. Science. 333:225-228.

[22] Godin, et al. (2010) Using buoyant mass to measure the growth of single cells. Nat Meth. 7:387-390.

[23] Pilizota T, Shaevitz JW (2012) Fast, multiphase volume adaptation to hyperosmotic shock by Escherichia coli. PLoS One. 7:e35205.

[24] Wood JM (1999) Osmosensing by Bacteria: Signals and Membrane-Based Sensors. Microbiol Mol Biol Rev. 62:230-262.

[25] McLaggan D (1994) Interdependence of K^{+} and glutamate accumulation during osmotic adaptation of Escherichia coli. 269:1911-1917.

[26] Milner JL (1988) Proline porter II is activated by a hyperosmotic shift in both whole cells and membrane vesicles of Escherichia coli K12. J Biol Chem. 263:14900-14905.

[27] Ray PM (1961) in Control Mechanisms in Cellular Processes, ed Bonner DM (Ronald Press, New York), pp 185-212.

[28] Proseus TE, Boyer JS (2008) Calcium pectate chemistry causes growth to be stored in Chara corallina: a test of the pectate cycle. Plant Cell Environ. 31:1147-1155.

[29] Levina N, et al. (1999) Protection of Escherichia coli cells against extreme turgor by activation of MscS and MscL mechanosensitive channels: identification of genes required for MscS activity. EMBO J. 18:1730-1737.

[30] Sukharev SI, Sigurdson WJ, Kung C, Sachs F (1999) Energetic and spatial parameters for gating of the bacterial large conductance mechanosensitive channel, MscL. J Gen Physiol. 113:525-539.

[31] Maja’s paper?

[32] Koch AL (1985) How bacteria grow and divide in spite of internal hydrostatic pressure. Can J Microbiol. 31:1071-1084.

[33] Lodish H, Berk A, Zipursky SL (2000) in Molecular Biology of the Cell, (W.H. Freeman, New York), Section 22.5.

[34] Deng Y, Sun M, Shaevitz J (2011) Direct measurement of cell wall stress stiffening and turgor pressure in live bacterial cells. Phys Rev Lett. 107:158101.

[35] Misra G, Rojas ER, Gopinathan A, Huang KC (2013) Mechanical consequences of cellwall turnover in the elongation of a Gram-positive bacterium. Biophys J. 104:2342.

[36] Schumann U, et al. (2010) YbdG in Escherichia coli is a threshold-setting mechanosensitive channel with MscM activity. Proc Natl Acad Sci USA. 107:12664-12669.

[37] Jaqaman K, et al. (2008) Robust single-particle tracking in live-cell time-lapse sequences. Nat Meth. 5:695-702.